Firms HL and LL are identical except for their financial leverage ratios and the interest rates they pay on debt. Each has $10 million in invested capital, has $1.5 million of EBIT, and is in the 40% federal-plus-state tax bracket. Firm HL, however, has a debt-to-capital ratio of 60% and pays 13% interest on its debt, whereas LL has a 20% debt-to-capital ratio and pays only 9% interest on its debt. Neither firm uses preferred stock in its capital structure.

  1. Calculate the return on invested capital (ROIC) for each firm. Round your answers to two decimal places.

    ROIC for firm LL is   %
    ROIC for firm HL is   %

  2. Calculate the rate of return on equity (ROE) for each firm. Round your answers to two decimal places.

    ROE for firm LL is    %
    ROE for firm HL is    %

  3. Observing that HL has a higher ROE, LL's treasurer is thinking of raising the debt-to-capital ratio from 20% to 60% even though that would increase LL's interest rate on all debt to 15%. Calculate the new ROE for LL. Round your answer to two decimal places.

    %
Answer #1

ROIC = Net operating profit after tax / Total capital

Net operating profit after tax = EBIT (1- Tax rate)

a) Total capital = $10 million

EBIT = $1.5 million

Tax rate = 40%

EBIT (1- Tax rate) = $1,500,000(1-0.4) = $900,000

ROIC= $900,000 / $10,000,000 = 0.09 = 9% for both LL and HL.

c) ROE= Net income / Shareholder's equity

HL: Net Income = EBIT(1- interest rate) (1- tax rate)

Net income= $1,500,000(1-0.13)(1-0.4) = $783,000

Total capital = $10million

Debt-to-capital ratio= 0.6

Hence equity is $4 million.

ROE = $783,000 / $4,000,000 = 0.19575 = 19.57%

d) LL: Net Income = EBIT(1- interest rate) (1- tax rate)

Net income= $1,500,000(1-0.09)(1-0.4) = $819,000

Total capital = $10million

Debt-to-capital ratio= 0.2

Hence equity is $8 million.

ROE = $819,000 / $8,000,000 = 0.1023575 = 10.24%

e) By plugging in the required interest rate and debt to capital ratio for LL:

LL: Net Income = EBIT(1- interest rate) (1- tax rate)

Net income= $1,500,000(1-0.15)(1-0.4) = $765,000

Total capital = $10million

Debt-to-capital ratio= 0.6

Hence equity is $4 million.

ROE = $765,000 / $4,000,000 = 0.19125 = 19.125% ~ 19.13%